OpenStudy (anonymous):

adult admission is 15.25, children is 4.50. during one day, a total of 2,800 adult and children were collected, and 28,725 in admission was collected. how many adults were collected.

OpenStudy (anonymous):

Are those 2800 adult + children admissions included in the 28725 admissions?

OpenStudy (anonymous):

$28,725 was collected

OpenStudy (anonymous):

Let x = number of adult tickets. Let y = number of child tickets. 15.25x + 4.50y = 28 725 x + y = 2800 Using the process of elimination: 15.25x + 4.50y = 28 725 4.50x + 4.50y = 12 600 = 10.75x = 16125 Therefore, x = 1500. Therefore, 1500 adult tickets were sold.